# Energy shifts and broadening of excitonic resonances in electrostatically-doped semiconductors

### v1

@@ I Introduction
Optical absorption and emission processes in electrostatically-doped semiconductors tell us a great deal about the pervading role of the Pauli exclusion principle. One ramification of the Pauli principle in semiconductors is the quantum-mechanical distinguishability of particles in excitonic complexes in terms of their valley and spin quantum numbers [1, 2, 3, 4, 5, 6, 7, 8]. In addition, the Pauli principle restricts the permissible types of recombination channels of dark excitons and trions [9, 10, 11, 12], the way hot excitons relax in energy [13, 14, 15, 16, 17, 18], and how impurities assist recombination processes [19]. However, when it comes to the relation between electrostatic doping and the broadening of excitonic optical resonances, or between electrostatic doping and the energy shifts of these resonances [20, 21, 22, 23, 24, 25, 26, 27], the Pauli exclusion principle is often given a lesser role and other mechanisms are invoked to explain such physics. To date, no proposed mechanism can satisfactorily or self-consistently explain the broadening and energy shifts of all optical resonances in charge-tunable semiconductors, including theoretical frameworks based on exciton-polarons [28, 29, 30, 31, 32], correlated trions (i.e. “tetrons”) [35, 33, 34, 36, 37, 38, 39], intervalley plasmons [40, 41, 42, 40], or arguments that rely on band filling and screening [43].
@@
To explain all experimental observations self-consistently and to inform on the essential role played by the Pauli exclusion principle, we propose in this work that all excitonic optical resonances should be classified as being either distinct or indistinct, and that their underlying excitonic complexes should be classified as being either optimal or suboptimal. We will show that this classification scheme provides a useful framework to understand whether and how various excitonic resonances shift and broaden in energy as the charge density is varied.
@@
Distinct versus indistinct optical resonances: When a semiconductor is doped with resident electrons in the conduction band (CB) or resident holes in the valence band (VB), there is an important difference between optical transitions that involve distinguishable versus indistinguishable e-h pairs. A distinguishable pair means that the electron and hole have unique quantum numbers that are not shared by any of the resident carriers. For example, when the electron and hole are photoexcited into unoccupied valleys [44]. Distinct (indistinct) optical resonances arise from optical transitions of distinguishable (indistinguishable) e-h pairs. To streamline the discussion in the rest of this work, we will also refer to resident carriers as indistinguishable if they share quantum numbers with the electron or hole in the pair, and as distinguishable otherwise.
@@
Optimal versus suboptimal excitonic complexes: We say that an excitonic complex is optimal if there are no resident carriers in the system with which the complex could bind to form a different lower-energy complex with more particles. Since particles with similar quantum numbers cannot be part of the same bound complex [45], the electrons (or holes) of an optimal excitonic complex are mutually distinguishable and they include all possible valley-spin species of resident electrons (or holes) at the edge of the CB (VB). This definition implies that the ground-state excitonic complex is always optimal. However, we will show that other complexes can be optimal as well.
@@
We will employ these definitions and present an overarching framework by which one can directly analyze the measured behavior of all optical resonances in charge-tunable semiconductors. Focusing on transition-metal dichalcogenide (TMD) monolayers and moiré heterobilayers by using the experimental findings of Refs. [7, 8, 11, 20, 21, 22, 23, 24, 25, 26, 27], we identify five important understandings and universal observations.
@@
    Distinct resonances of optimal excitonic complexes neither broaden nor decay when resident carriers are added to the semiconductor. Examples include certain types of trions in valley-polarized monolayers, 6-particle hexcitons, or intralayer excitons in moiré heterobilayers. These resonances only show a steady energy redshift that is likely caused by a small difference between renormalizations of the band gap and binding energies. There is no broadening or decay in this case because the absorption or recombination processes do not conflict with Pauli exclusion and therefore there is no need to shake-up the Fermi sea (i.e., to rearrange the resident carriers).
@@
    With increasing charge density, the energy blueshift and broadening of the neutral-exciton absorption resonance are especially strong in the presence of distinguishable resident carriers. Their shakeup leads to a stronger blueshift than the blueshift caused by Pauli blocking (band filling). That is, stronger in comparison to the increase of the optical energy gap when indistinguishable resident carriers occupy low energy states in the photoexcited valley.
@@
Before embarking, it is important to state that our analysis is valid in cases that the neutral exciton binding energy is much larger than the Fermi energy or cyclotron energy when applying a magnetic field, $\varepsilon_{X}\gg E_{F},\,\hbar\omega_{c}$, meaning that the effective Bohr radius of the exciton is smaller than the cyclotron radius or the average distance between resident carriers.
@@
Section II outlines the classification of optical resonances and excitonic complexes in MoSe2 and WSe2 monolayers based on their distinctiveness and optimality. Section III discusses various experimental results and then focuses on the universal energy redshift of distinct resonances of optimal excitonic complexes when the charge density increases in monolayers and moiré heterobilayers. We then show that screening by resident carriers is not leading to the decay of excitonic complexes and that band filling is not the primary source of energy blueshift. Section IV deals with shakeup processes, first focusing on the broadening and energy shifts of indistinct resonances of optimal excitonic complexes, and then focusing on the energy blueshift of photoexcited neutral excitons under various conditions. Section V deals with the relation between the compressibility of the resident carrier gas and the behavior of neutral exciton resonances in monolayers and moiré heterobilayers. A summary is given in Sec. VI.
@@
Figures 1(a) and (b) show the photoexcitation in electron-rich conditions. Both exciton and trion resonances are said to be indistinct because the photoexcited electron is promoted to a CB that is already occupied by resident electrons having the same quantum numbers (spin and valley). The neutral exciton in Fig. 1(a) is suboptimal because there are opposite-spin electrons in the $-K$ CB with which this exciton could bind to form a lower-energy complex with more particles (i.e., a trion). Conversely, the negative trion shown in Fig. 1(b) is optimal because the photoexcited pair binds to an electron from the valley at $-K$, and there are no other kinds of resident electrons with which the pair can bind. The trion is depicted as the binding between the photoexcited e-h pair and particle-hole excitation of the Fermi sea in the $-K$ valley, and the entire complex is known as a tetron, wherein the trion moves together and is correlated with the CB hole in the Fermi sea around the trion [35, 33, 34, 36]. Figures 1(c) and (d) show the respective physics in a hole-rich MoSe2 monolayer.
Figure 1: Photoexcitation of the $K$ valley in an electrostatically-doped MoSe2 monolayer. Valleys with solid (dashed) lines host spin-up (spin-down) electrons. The diagrams describe neutral excitons ($X^{0}$) and trions ($X^{\pm}$). For the latter, the photoexcited e-h pair binds to a charge particle from the $-K$ valley, shown by a particle-hole excitation of the Fermi sea. The diagrams in (e)-(f) describe photoexcitation under a strong magnetic field $B$, such that Landau quantization of VB holes (CB electrons) happens solely at the $K$ ($-K$) valley. All optical transitions except for (f) are indistinct resonances, because the photoexcited electron or hole is injected into already-occupied bands.
@@
The excitonic complexes in WSe2 (or WS2) monolayers are similar in character to those of MoSe2 for the case of hole doping, but are qualitatively different for the case of electron doping. This is because the sign of the CB spin-orbit splitting is opposite in W-based TMD monolayers in comparison to MoSe2. As shown in Fig. 2, photoexcited electrons are promoted to the upper CBs, rendering the photoexcited e-h pair distinguishable in electron-rich conditions as long as the upper CBs are not populated by resident electrons. The photoexcitation involves distinct resonances for the neutral exciton in Fig. 2(a), singlet trion in (b), triplet trion in (c), and the optimal hexciton complex in (d). The hexciton comprises six particles: the photoexcited e-h pair, and two particle-hole pairs from the two lower CBs at ±$\pm K$ [25, 24, 39] (Appendix A). When the upper CB valleys start to fill at high doping levels, the photoexcited e-h pair can form an eight-particle oxciton by binding to an additional particle-hole pair [24], as shown in Fig. 2(e). The oxciton becomes the optimal complex but it has indistinct resonance at $B$=0. A strong magnetic field can turn this resonance distinct again by emptying the upper CB at $K$, as shown in Fig. 2(f).
@@ III.1 Universal energy redshift of distinct resonances of optimal excitonic complexes
We first focus on the energy redshift of distinct resonances that arise from optimal excitonic complexes. At zero magnetic field, these complexes are type-A hexcitons in electron-rich W-based monolayers and type-B hexcitons in any hole-rich TMD monolayer. Their resonances are marked by arrows in Figs. 4(a) and (b), and their transition diagrams are shown in Figs. 2(d), 3(b), and 3(c). Figure 4(a) shows that the type-A hexciton resonance in WSe2 remains sharp and exhibits a pronounced redshift with increasing electron density, which abruptly ceases only when the gate voltage reaches $\sim$11 V (electron density $\sim$5$\times$1012 cm-2) and electrons begin filling the upper CBs [20]. Beyond this point the photoexcited e-h pair becomes indistinguishable from some carriers in the Fermi sea, and the resonance therefore becomes indistinct. On the other hand, type-B hexcitons in hole-doped conditions remain optimal and their resonances remain distinct even at very large hole densities because $\Delta_{v}\gg E_{F}$, as depicted in Figs. 3(b) and (c). These resonances, labeled by $H_{\mathcal{B}}$ in Figs. 4(a) and (b), exhibit a robust redshift and do not broaden even up to the highest reported doping. We note that type-B resonances are typically broader than those of type-A, due to their very short lifetime [47].
@@
Starting with MoSe2 in Fig. 4(d), the $X^{\pm}$ resonances in $B$=16 T exhibit a pronounced redshift with increasing doping (in marked contrast to their negligible redshift at $B$=0). However, this redshift persists only as long as the optical transition involves a distinguishable e-h pair. The redshift of $X^{-}$ starts from its emergence when the monolayer becomes electron-doped at $V_{g}\sim$6 V and stops at $V_{g}\sim$10 V (indicated by the white dotted line near $X^{-}$ in Fig. 4(d)). Beyond this point, the Fermi energy exceeds the Zeeman energy and the photoexcited valley starts to fill, and the $X^{-}$ resonance is no longer distinct, marking the onset of its broadening and energy blueshift. Referring to the diagram of Fig. 1(f), the redshift ceases and the broadening is enhanced when the electron density is large enough to begin filling the CB valley at $K$.
@@
Finally, we focus on the electron-doped side of Fig. 4(c), where the gate voltage window of fully valley-polarized electrons is narrowest in W-based monolayers because of the smallness of $g_{c,1}$. The emergence of hexciton resonances ($H$) and the concomitant decay of trions ($X^{-}_{S,T}$) takes place when LLs are filled in the lower CB valleys of both ±$\pm K$ [24]. Similar to the case at $B=0$ in Fig. 4(a), the energy redshift of the type-A hexciton resonance persists only as long as the upper CB valleys are empty (i.e., as long as it remains a distinct resonance). The crossover from hexciton to oxciton occurs when the upper CB valley at $-K$ starts to fill, i.e. when $V_{g}\gtrsim 3.5$ V in this device (electron densities $\gtrsim$5$\times$1012 cm-2). The diagram of this distinct and optimal oxciton is shown in Fig. 2(f). The observed $H\,\rightarrow\,O$ crossover is highlighted in the inset of Fig. 4(c), which magnifies the reflectance signal with helicity $\sigma_{+}$ inside the dashed box at the top left corner (Appendix B).
@@ III.2 Suggested origin of the universal energy redshift
These data show that distinct resonances of optimal excitonic complexes exhibit a universal energy redshift when the charge density increases, and further, that these resonances neither decay nor broaden. The lack of decay and broadening can be reasoned by noting that a distinguishable e-h pair can bind to resident carriers with all available quantum numbers without having to perturb the distribution of resident carriers that are not part of the optimal complex (Appendix A). In Sec. IV, we will discuss why and how shakeup processes broaden the optical resonances in cases that the e-h pair is indistinguishable and/or when the complex is suboptimal.
The universal energy redshift is common to distinct resonances of various optimal complexes. This commonality can be attributed to the small but non-zero difference that exists between band-gap renormalization (BGR) and reduction of the binding energy due to screening by resident carriers (namely, $\Delta E=\Delta E_{g}-\Delta E_{b}$). Increasing the charge density from $n_{1}$ to $n_{2}=n_{1}+\delta n$ is equivalent to increasing the screening through the effective dielectric constant from $\epsilon_{1}$ to $\epsilon_{2}=\epsilon_{1}+\delta\epsilon$. The BGR of a semiconductor is attributed to the Coulomb hole effect [50], reflecting the change in energy needed to excite an electron across the band gap when the Coulomb potential at the immediate vicinity of the electron is changed. Considering a small change, such that $\delta\epsilon$ is small and positive, we get [50, 51, 52, 53]
@@
For simplicity, the derivation above was made for a neutral exciton. Yet, the redshift effect should be similar for complexes with larger number of particles because most of the binding energy comes from the photoexcited pair, where additional particles contribute less and less energy. For example, if we consider a WSe2 monolayer that is encapsulated in hexagonal boron nitride, then the gained energy by forming an exciton is about 170 meV compared with an unbound electron-hole pair [55]. Adding an electron to form a trion gains extra $\sim$30 meV [1], adding the next electron to form a hexciton gains extra $\sim$10 meV, and the next electron to form an oxciton gains extra 3 meV [25, 24]. This behavior is similar to that of energy levels of successive electron shells in atoms but with smaller energy scales.

### v2

@@ I Introduction
Optical absorption and emission processes in electrostatically-doped semiconductors tell us a great deal about the pervading role of the Pauli exclusion principle. One ramification of the Pauli principle is that under most conditions, excitonic complexes in semiconductors are composed of particles with different valley and spin quantum numbers [1, 2, 3, 4, 5, 6, 7, 8, 9]. In addition, the Pauli principle restricts the permissible types of recombination channels of dark excitons and trions [10, 11, 12, 13], the way hot excitons relax in energy [14, 15, 16, 17, 18, 19], and how impurities assist recombination processes [20]. However, when it comes to the relation between electrostatic doping and the broadening of excitonic optical resonances, or between electrostatic doping and the energy shifts of these resonances [21, 22, 23, 24, 25, 26, 27, 28], the Pauli exclusion principle is often given a lesser role and other mechanisms are invoked to explain such physics. To date, no proposed mechanism can satisfactorily or self-consistently explain the broadening and energy shifts of all optical resonances in charge-tunable semiconductors, including theoretical frameworks based on exciton-polarons [29, 30, 31, 32, 33], correlated trions (i.e. “tetrons”) [35, 34, 36, 37, 38, 39], intervalley plasmons [40, 41, 42, 40], or arguments that rely on band filling and screening [43, 44, 45, 46].
@@
To explain all experimental observations self-consistently and to inform on the essential role played by the Pauli exclusion principle, we propose in this work that all excitonic optical resonances should be classified as being either distinct or indistinct, and that their underlying excitonic complexes should be classified as being either optimal or suboptimal. We first provide formal definitions for the distinctiveness of resonances and the optimality of excitonic complexes. In the next sections and with the aid of the examples shown in Figs. 1-3, we will show how to apply these definitions and how this classification scheme provides a useful framework to understand whether and how various excitonic resonances shift and broaden in energy as the charge density is varied.
@@
Distinct versus indistinct optical resonances: When a semiconductor is doped with resident electrons in the conduction band (CB) or resident holes in the valence band (VB), there is an important difference between optical transitions that involve distinguishable versus indistinguishable e-h pairs. A distinguishable pair means that the electron and hole have unique quantum numbers that are not shared by any of the resident carriers. For example, when the electron and hole are photoexcited into unoccupied valleys. Distinct (indistinct) optical resonances arise from optical transitions of distinguishable (indistinguishable) e-h pairs. To streamline the discussion in the rest of this work, we will also refer to resident carriers as indistinguishable if they share quantum numbers with the electron or hole in the pair, and as distinguishable otherwise.
@@
Optimal versus suboptimal excitonic complexes: We say that an excitonic complex is optimal if there are no resident carriers in the system with which the complex could bind to form a different lower-energy complex with more particles. Other than in very special circumstances (Appendix A), particles with similar quantum numbers cannot be part of the same bound complex [47, 48, 49]. As such, the electrons (or holes) of an optimal excitonic complex are mutually distinguishable and they include all possible valley-spin species of resident electrons (or holes) at the edge of the CB (VB). This definition implies that the ground-state excitonic complex is always optimal. However, we will show that other complexes can be optimal as well.
@@
We will employ these definitions and present an overarching framework by which one can directly analyze the measured behavior of all optical resonances in charge-tunable semiconductors. Focusing on transition-metal dichalcogenide (TMD) monolayers, their moiré heterobilayers, and conventional III-V and II-VI semiconductor quantum wells (QWs) by using the experimental findings of Refs. [7, 8, 12, 21, 22, 23, 24, 25, 26, 27, 28, 50, 51, 52], we identify five important understandings and universal observations.
@@
    Distinct resonances of optimal excitonic complexes neither broaden nor decay when resident carriers are added to the semiconductor. Examples include certain types of trions in valley- or spin-polarized semiconductors, 6-particle hexcitons, or intralayer excitons in moiré heterobilayers. These resonances only show a steady energy redshift that is likely caused by a small difference between renormalizations of the band gap and binding energies. There is no broadening or decay in this case because the absorption or recombination processes do not conflict with Pauli exclusion and therefore there is no need to shake-up the Fermi sea (i.e., to rearrange the resident carriers).
@@
    With increasing charge density, the energy blueshift and broadening of the neutral-exciton absorption resonance are especially strong in the presence of distinguishable resident carriers. The energy blueshift stems from transferring the spectral weight of the suboptimal exciton to the optimal trion resonance. The resulting energy blueshift is stronger than the Moss–Burstein effect [53, 54], which is caused by Pauli blocking (band filling) when indistinguishable resident carriers occupy low energy states in the photoexcited valley.
@@
Before embarking, it is important to state that our analysis is valid in cases where the neutral exciton binding energy is much larger than the Fermi energy or cyclotron energy when applying a magnetic field, $\varepsilon_{X}\gg E_{F},\,\hbar\omega_{c}$, meaning that the effective Bohr radius of the exciton is smaller than the cyclotron radius or the average distance between resident carriers.
@@
Section II outlines the classification of optical resonances and excitonic complexes in MoSe2 and WSe2 monolayers based on their distinctiveness and optimality. Section III discusses various experimental results and then focuses on the universal energy redshift of distinct resonances of optimal excitonic complexes when the charge density increases in TMD monolayers, moiré heterobilayers, and conventional semiconductor QWs. We then show that screening by resident carriers is not leading to the decay of excitonic complexes. Section IV deals with shakeup processes that correspond to the broadening and energy shifts of indistinct resonances of optimal excitonic complexes. Section V analyzes the energy blueshift of photoexcited neutral excitons under various absorption conditions. Section VI deals with the relation between the compressibility of the resident carrier gas and the behavior of neutral exciton resonances in monolayers and moiré heterobilayers. A summary is given in Sec. VII. Appendices A-J elaborate on key conclusions made in the main text and they provide technical details of calculated results.
@@
Figure 1: Photoexcitation of the $K$ valley in an electrostatically-doped MoSe2 monolayer. Valleys with solid (dashed) lines host spin-up (spin-down) electrons. The diagrams describe neutral excitons ($X^{0}$) and trions ($X^{\pm}$). For the latter, the photoexcited e-h pair binds to a charge particle from the $-K$ valley, shown as a particle-hole pair of the Fermi sea. The diagrams in (e)-(f) describe photoexcitation under a strong magnetic field $B$, such that Landau quantization of VB holes (CB electrons) happens solely at the $K$ ($-K$) valley. All optical transitions except for (f) are indistinct resonances, because the photoexcited electron or hole is injected into already-occupied bands.
Figures 1(a) and (b) show the photoexcitation in electron-rich conditions. Both exciton and trion resonances are said to be indistinct because the photoexcited electron is promoted to a CB that is already occupied by resident electrons having the same quantum numbers (spin and valley). Moreover, the neutral exciton in Fig. 1(a) is suboptimal because there are opposite-spin electrons in the $-K$ CB with which this exciton could bind to form a lower-energy complex with more particles (i.e., a trion). Conversely, the negative trion shown in Fig. 1(b) is optimal because the photoexcited pair binds to an electron from the valley at $-K$, and there are no other kinds of resident electrons with which the pair can bind. The trion is depicted as the binding between the photoexcited e-h pair and particle-hole pair from the Fermi sea in the $-K$ valley, and the entire complex is known as a tetron, wherein the trion moves together and is correlated with the CB hole in the Fermi sea around the trion [35, 34, 36]. Figures 1(c) and (d) show the corresponding transitions in a hole-rich MoSe2 monolayer.
@@
The excitonic complexes in WSe2 (or WS2) monolayers are similar in character to those of MoSe2 for the case of hole doping, but are qualitatively different for the case of electron doping. This is because the sign of the CB spin-orbit splitting is opposite in W-based TMD monolayers in comparison to MoSe2. As shown in Fig. 2, photoexcited electrons are promoted to the upper CBs, rendering the photoexcited e-h pair distinguishable in electron-rich conditions as long as the upper CBs are not populated by resident electrons. The photoexcitation involves distinct resonances for the neutral exciton in Fig. 2(a), singlet trion in (b), triplet trion in (c), and the optimal hexciton complex in (d). The hexciton comprises six particles: the photoexcited e-h pair, and two particle-hole pairs from the two lower CBs at ±$\pm K$ [26, 25, 39] (Appendix B). When the upper CB valleys start to fill at high doping levels, the photoexcited e-h pair can form an eight-particle oxciton by binding to an additional particle-hole pair [25], as shown in Fig. 2(e). The oxciton becomes the optimal complex but it has an indistinct resonance at $B$=0. A strong magnetic field can turn this resonance distinct again by emptying the upper CB at $K$, as shown in Fig. 2(f).
@@ III.1 Universal energy redshift of distinct resonances of optimal excitonic complexes
We first focus on the energy redshift of distinct resonances that arise from optimal excitonic complexes. At zero magnetic field, these complexes are type-A hexcitons in electron-rich W-based monolayers and type-B hexcitons in any hole-rich TMD monolayer. Their resonances are marked by arrows in Figs. 4(a) and (b), and their transition diagrams are shown in Figs. 2(d), 3(b), and 3(c). Figure 4(a) shows that the type-A hexciton resonance in WSe2 remains sharp and exhibits a pronounced redshift with increasing electron density, which abruptly ceases only when the gate voltage reaches $\sim$11 V (electron density $\sim$6$\times$1012 cm-2) and electrons begin filling the upper CBs [21]. Beyond this point the photoexcited e-h pair becomes indistinguishable from some carriers in the Fermi sea, and the resonance therefore becomes indistinct. On the other hand, type-B hexcitons in hole-doped conditions remain optimal and their resonances remain distinct even at very large hole densities because $\Delta_{v}\gg E_{F}$, as depicted in Figs. 3(b) and (c). These resonances, labeled by $H_{\mathcal{B}}$ in Figs. 4(a) and (b), exhibit a robust redshift and do not broaden even up to the highest reported doping. We note that type-B resonances are typically broader than those of type-A, due to their very short lifetime [56].
@@
Starting with MoSe2 in Fig. 4(d), the $X^{\pm}$ resonances in $B$=16 T exhibit a pronounced redshift with increasing doping (in marked contrast to their negligible redshift at $B$=0). However, this redshift persists only as long as the optical transition involves a distinguishable e-h pair. The redshift of $X^{-}$ starts from its emergence when the monolayer becomes electron-doped at $V_{g}\sim$6 V and stops at $V_{g}\sim$10 V (indicated by the white dotted line near $X^{-}$ in Fig. 4(d)). Beyond this point, the Fermi energy exceeds the Zeeman energy and the photoexcited valley starts to fill, and the $X^{-}$ resonance is no longer distinct, marking the onset of its broadening and energy blueshift. This behavior is consistent with other observations seen at different magnetic fields and electron densities [57]. Referring to the diagram of Fig. 1(f), the redshift ceases and the broadening is enhanced when the electron density is large enough to begin filling the CB valley at $K$.
@@
Finally, we focus on the electron-doped side of Fig. 4(c), where the gate voltage window of fully valley-polarized electrons is narrowest in W-based monolayers because of the smallness of $g_{c,1}$. The emergence of hexciton resonances ($H$) and the concomitant decay of trions ($X^{-}_{S,T}$) takes place when LLs are filled in the lower CB valleys of both ±$\pm K$ [25]. Similar to the case at $B=0$ in Fig. 4(a), the energy redshift of the type-A hexciton resonance persists only as long as the upper CB valleys are empty (i.e., as long as it remains a distinct resonance). The crossover from hexciton to oxciton occurs when the upper CB valley at $-K$ starts to fill, i.e. when $V_{g}\gtrsim 3.5$ V in this device (electron densities $\gtrsim$6$\times$1012 cm-2). The diagram of this distinct and optimal oxciton is shown in Fig. 2(f). The observed $H\,\rightarrow\,O$ crossover is highlighted in the inset of Fig. 4(c), which magnifies the reflectance signal with helicity $\sigma_{+}$ inside the dashed box at the top left corner (Appendix C).
@@ III.2 Suggested origin of the universal energy redshift
These data show that distinct resonances of optimal excitonic complexes exhibit a universal energy redshift when the charge density increases, and further, that these resonances neither decay nor broaden. The lack of decay and broadening can be reasoned by noting that a distinguishable e-h pair can bind to resident carriers with all available quantum numbers without having to perturb the distribution of resident carriers that are not part of the optimal complex (Appendix B). In Sec. IV, we will discuss why and how shakeup processes broaden the optical resonances in cases that the e-h pair is indistinguishable.
The universal energy redshift is common to distinct resonances of various optimal complexes. This commonality can be attributed to the small but non-zero difference that exists between band-gap renormalization (BGR) and reduction of the binding energy due to screening by resident carriers (namely, $\Delta E=\Delta E_{g}-\Delta E_{b}$). Increasing the charge density from $n_{1}$ to $n_{2}=n_{1}+\delta n$ is equivalent to increasing the screening through the effective dielectric constant from $\epsilon_{1}$ to $\epsilon_{2}=\epsilon_{1}+\delta\epsilon$. The BGR of a semiconductor is attributed to the change in energy needed to excite an electron across the band gap when the Coulomb potential at the immediate vicinity of the electron is changed [60]. Considering a small change, such that $\delta\epsilon$ is small and positive, we get [61, 62, 63, 64, 60]
@@
For simplicity, the derivation above was made for a neutral exciton. Yet, the energy redshift of complexes with larger number of particles should be similar for two reasons. First, the energy shift that one observes in experiment is a measure of the change to the optical gap of the photoexcited (or recombined) electron-hole pair, whereas changes in the self-energy of other particles in the complex are not probed in these optical experiments. For example, a trion inherits the self-energy of a resident carrier to which the e-h pair binds, and therefore, the absorbed photon energy cannot inform us on a self-energy that remains the same before and after the photoexcitation. More on this subtle point is discussed in Appendix D. The second reason that the redshift is largely independent of the number of particles in the complex is that most of the binding energy comes from the photoexcited pair, where adding particles contributes less and less energy. For example, if we consider a WSe2 monolayer that is encapsulated in hexagonal boron nitride, then the gained energy by forming an exciton is about 170 meV compared with an unbound electron-hole pair [66]. Adding an electron to form a trion gains extra $\sim$30 meV [1], adding the next electron to form a hexciton gains extra $\sim$10 meV, and the next electron to form an oxciton gains extra 3 meV [26, 25]. This behavior is similar to that of energy levels of successive electron shells in atoms but with smaller energy scales.